# FORCE: Transferable Visual Jailbreaking Attacks via Feature Over-Reliance CorrEction

### v2

@@ Abstract
Warning: This paper contains content that may be considered harmful.
The integration of new modalities enhances the capabilities of multimodal large language models (MLLMs) but also introduces additional vulnerabilities. In particular, simple visual jailbreaking attacks can manipulate open-source MLLMs more readily than sophisticated textual attacks. However, these underdeveloped attacks exhibit extremely limited cross-model transferability, failing to reliably identify vulnerabilities in closed-source MLLMs. In this work, we analyse the loss landscape of these jailbreaking attacks and find that the generated attacks tend to reside in high-sharpness regions, whose effectiveness is highly sensitive to even minor parameter changes during transfer. To further explain the high-sharpness localisations, we analyse their feature representations in both the intermediate layers and the spectral domain, revealing an improper reliance on narrow layer representations and semantically poor frequency components. Building on this, we propose a Feature Over-Reliance CorrEction (FORCE) method, which guides the attack to explore broader feasible regions across layer features and rescales the influence of frequency features according to their semantic content. By eliminating non-generalizable reliance on both layer and spectral features, our method discovers flattened feasible regions for visual jailbreaking attacks, thereby improving cross-model transferability. Extensive experiments demonstrate that our approach effectively facilitates visual red-teaming evaluations against closed-source MLLMs.
@@ 1 Introduction
To meet the growing demand for complex tasks, the capability to process multimodal information has been rapidly integrated into multimodal large language models (MLLMs) (OpenAI, 2025; Anthropic, 2025; Google, 2025). Despite their remarkable performance, the increasing deployment of these models in decision-critical domains has raised societal concerns about their potential risks (Perez et al., 2022; Ganguli et al., 2022). Recent red-teaming efforts reveal that, although MLLMs exhibit strong safeguards against textual jailbreaking attacks, they can be easily manipulated through vulnerabilities introduced by newly embedded modalities (Qi et al., 2024; Bailey et al., 2023).
Among various attacks, optimisation-based visual jailbreaking attacks are considered one of the most effective for identifying vulnerabilities in MLLMs, as they can reliably bypass the safety guardrails of open-source models with imperceptible perturbations (Zhao et al., 2023; Niu et al., 2024; Aichberger et al., ). As illustrated in Figure 1, visual attacks optimised on the source model can effectively exploit its inherent vulnerabilities and elicit harmful responses to malicious instructions, whereas the same requests are refused when paired with a non-adversarial image. Nevertheless, these visual attacks exhibit extremely limited cross-model transferability (Schaeffer et al., 2025), as the exploited vulnerabilities are specific to the source MLLM and fail to generalise to target MLLMs during transfer. Consequently, such attacks fall short of posing a practical threat to closed-source commercial MLLMs and remain inadequate for real-world red-teaming evaluations.
@@
Motivated by this, we further analyse the feature representations of visual jailbreaking attacks in both the intermediate layers and the spectral domain, uncovering the existence of non-generalizable reliance. Specifically, the feasible regions of visual attacks display distinct characteristics across layers. Closer to the earlier layers, these attacks depend more heavily on model-specific features to mislead MLLMs, resulting in narrower and more fragile feasible regions. Regarding the spectral domain, we observe that as optimisation progresses, high-frequency information exerts increasing influence on attack effectiveness, eventually surpassing low-frequency components that contain richer semantic content. This trend suggests an overemphasis on high-frequency information, making the generated attacks depend on semantically weak features that lack generalisability. Both aspects of improper feature reliance hinder visual jailbreaking attacks from capturing robust representations, which in turn confines them to high-sharpness regions and results in poor cross-model transferability.
@@ 2 Related Work
Multimodal Large Language Models. There are two mainstream architectures for integrating new modalities: adapter-based MLLMs (Liu et al., 2023b; Zhu et al., 2023; Bai et al., 2023a) and early-fusion MLLMs (Zhou et al., 2024; Xiao et al., 2024; Team, 2024). Adapter-based MLLMs employ an adapter to project the output of an image encoder, such as CLIP (Radford et al., 2021), into the embedding space of the large language models (LLMs). On the other hand, early-fusion MLLMs utilise a unified tokeniser to process multimodal information within a shared embedding space. Both designs can leverage the powerful reasoning and understanding capabilities of LLMs to support a wide range of multimodal tasks, with outputs predicted according to the joint conditional distribution of textual and visual information, $p_{\bm{\theta}}\left(\mathbf{y}\mid\mathbf{x}_{\text{img}},\mathbf{x}_{\text{txt}}\right)$.
@@
Textual Jailbreaking Attack. Jailbreaking attacks arise from the discovery that hand-crafted adversarial prompts can bypass safeguards in LLMs, leading them to answer malicious queries and produce harmful content (Shen et al., 2023; Liu et al., 2023c). To automatically uncover vulnerabilities in LLMs, three types of jailbreaking attack strategies have been rapidly developed. Heuristic-based attacks typically leverage genetic algorithms to modify a prototype corpus until they successfully bypass the safety guardrails (Liu et al., ; Shah et al., 2023; Yu et al., 2023). LLM-based attacks utilise the inherent capabilities of LLMs to rewrite malicious queries, obstructing the victim model’s perception (Chao et al., 2023; Yao et al., 2023). Optimisation-based attacks define an affirmative target output and leverage gradient information to iteratively update the adversarial suffix, ultimately eliciting undesirable responses (Zou et al., 2023; Yang et al., 2025a; Liao & Sun, 2024).
@@
Although the aforementioned textual attacks can also manipulate MLLMs, their effectiveness diminishes with the growing strength of textual alignment (Touvron et al., 2023; Bai et al., 2022; Rafailov et al., 2024). In contrast, MLLMs demonstrate relatively weak alignment regarding vulnerabilities associated with new modalities (Shayegani et al., 2023; Schaeffer et al., 2025), thereby establishing visual jailbreaking attacks as a promising direction for red-teaming evaluations.
@@
Visual Jailbreaking Attack. Visual jailbreaks are typically classified into two categories: generation-based and optimisation-based methods. Generation-based methods either craft image typography to encode malicious textual content (Li et al., 2024; Yang et al., 2025b) or generate harmful images matching the textual semantics (Teng et al., 2024; Zhao et al., 2025). These generated malicious images can mislead MLLMs through the visual modality while simultaneously circumventing textual alignment mechanisms. However, such methods depend on human effort or auxiliary models to produce required visual typography or query–image pairs, making them resource-intensive. More importantly, this type of method lacks the ability to capture the fine-grained vulnerabilities, falling short of reliably manipulating the MLLMs (Schaeffer et al., 2025).
@@
In contrast, optimisation-based methods, such as the Projected Gradient Descent (PGD) attack (Madry et al., 2018) and its variants (Zhao et al., 2023; Qi et al., 2024; Bailey et al., 2023; Niu et al., 2024), use gradient information to optimise the jailbreaking perturbation $\delta$, thereby reliably exposing model vulnerabilities. In these methods, an affirmative target output of length $S$, such as Sure, here is, is first defined, and then the loss is calculated as:
@@
Despite achieving near-perfect success in manipulating the source MLLM, optimisation-based methods generate visual attacks with limited transferability to target MLLMs (Schaeffer et al., 2025). To thoroughly assess and expose potential risks in closed-source LLMs, this work aims to understand and improve the transferability of optimisation-based visual jailbreaking attacks.
@@ 3.1 Loss Landscape of Visual Jailbreaking Attack
As shown in Figure 1, while optimisation-based visual jailbreaking attacks can easily bypass the safety guardrails of victim MLLMs, their limited transferability to target models constrains their real-world practicality. Inspired by prior research on classification tasks (Chen et al., 2023; Wei et al., 2023), we first investigate the transferability of visual jailbreaking attacks through the geometry of the loss landscape. Throughout this section, we use LLaVA-v1.5-7B (Liu et al., 2023a) as the source MLLM, adopt standard PGD (Madry et al., 2018) with a step size of $2/255$ and a perturbation budget of $32/255$, and set ‘‘Sure, here is’’ as the optimisation target.
Figure 2: The input (top) and weight (bottom) loss landscape of the visual jailbreaking attack. The blue and yellow points correspond to successful and failed examples on the source MLLM, respectively.
First, we visualise the input loss landscapes of visual jailbreaking attacks by introducing pixel perturbations in two directions, one aligned with the gradient and the other randomly sampled from a uniform distribution. As observed in Figure 2 (top), the generated visual attacks effectively manipulate the source MLLM to achieve the optimisation objective, as evidenced by the nearly 0 loss at the original point. However, when we inject small pixel perturbations, the loss increases sharply, reflecting that the attack rapidly loses its effectiveness in misleading the model. For instance, even a 0.03 pixel perturbation along the adversarial direction can raise the loss above 0.28, which is sufficient to invalidate the attack. We also introduce weight perturbations to the model parameters to simulate the impact of transfer-induced parameter shifts on attack effectiveness. As depicted in Figure 2 (bottom), we observe that the attack is trapped in a local optimum of the source MLLM, where even a minor weight perturbation of 0.0002 can push it out of the feasible region and render it ineffective. This sharp loss landscape indicates that optimisation-based methods tend to rely on model-specific features, which are sensitive to minor changes and result in unreliable performance when generalised to target models.
@@ 3.2 Features Representations on Different Layers
To disentangle the feature reliance responsible for high-sharpness regions, we conduct a detailed analysis of the intermediate layer representations of generated visual attacks. For a fair comparison, we separately extract each layer’s features from a successful visual jailbreaking attack and a natural image, and then construct interpolated representations using the convex combination $(1-\mu)\cdot f_{\theta}(\text{jail})+\mu\cdot f_{\theta}(\text{nat})$, to exclude inter-layer differences such as parameter norms and activation scales. We also interpolate features between two different visual jailbreaking attacks, as detailed in Appendix A.
@@
However, toward shallower layers, visual jailbreaking attacks exhibit progressively narrower feasible regions in the feature space. As evidenced by Figure 3, in the 11th layer, the attack must retain more than 90% of adversarial features to successfully manipulate the source MLLM, while the introduction of merely 30% of natural features is sufficient to drive the loss sharply beyond 1.2. These observations suggest that in shallower layers, visual jailbreaking attacks exhibit an increasing reliance on model-specific features, manifested as narrower feasible regions. This reliance on non-generalizable early-layer features, in turn, confines the generated attacks to high-sharpness regions of the input space, making them unstable when transferred to other models.
@@ 3.3 Influence of Different Frequency Features
In addition to layer-wise features, we also examine the role of spectral information in visual jailbreaking attacks during the optimisation. Specifically, we first apply a Fourier transform to the visual attack and divide the spectrum into ten equal-width frequency bands (Kim et al., 2024). Then, we independently mask each frequency band and reconstruct the image via inverse Fourier transform. Finally, we compute the loss of the masked attacks to evaluate their reliance on spectral features.
@@
Nevertheless, as optimisation proceeds, the attack’s effectiveness becomes increasingly dependent on high-frequency components. As shown in Figure 4, at the 350th iteration, the 50–60% and 60–70% spectral features exhibit a more pronounced influence than at the 250th iteration, and removing them causes a greater degradation in attack effectiveness than the lower-frequency 40–50% range. This anomalous trend intensifies with further optimisation. By the 750th iteration, removing the third-highest frequency band alone is sufficient to make the visual jailbreaking attack fail to mislead the source MLLM. This trend indicates that visual jailbreaking attacks tend to increasingly rely on high-frequency features to mislead MLLMs, grounding their success in superficial patterns rather than semantically meaningful content. Such overemphasis on non-generalizable features makes the generated attacks highly model-specific and undermines their transferability across different MLLMs.
@@
3.4 Feature Over-Reliance CorrEction Method
$d_{l}=\|\left(f_{\theta,l}(\mathbf{x}_{\text{img}}+\delta,\mathbf{x}_{\text{txt}})\right)-\left(f_{\theta,l}(\mathbf{x}_{\text{img}}+\delta+\eta,\mathbf{x}_{\text{txt}})\right)\|_{2}^{2},\quad l=1,\dots,L.$ (3)
To identify the spectral features with excessive influence, we separately mask $M$ equal-width frequency bands $B_{m}$, and calculate their associated losses $\ell_{m}$, similar to Section 3.3. To restore the natural distribution, where semantic content plays a principal role in model perception, we downscale high-frequency components only when they exert greater influence than their corresponding low-frequency counterparts:
$w_{m}$ $=\min\!\left(1,\tfrac{\ell_{m-1}}{\ell_{m}}\right),\quad m=1,\ldots,M.$ (7)
$S$ $=\sum\nolimits_{m=1}^{M}(w_{m}\cdot\mathbbm{1}_{B_{m}}).$
Subsequently, we perform an element-wise multiplication of the frequency scaling matrix $S$ with the magnitude spectrum $A$ obtained from the Fourier transform $(A,\Phi)\leftarrow\mathrm{FFT}(\delta)$, and reconstruct the jailbreaking perturbation via the inverse Fourier transform $\delta_{\text{rescaled}}=\mathrm{IFFT}\left((A\odot S)\odot e^{\mathrm{i}\Phi}\right)$. We integrate these two components into a standard PGD algorithm by first rescaling the abnormal frequency bands and then exploring broader layer representations. This design eliminates non-generalizable feature reliance and encourages a flatter loss landscape for the generated visual jailbreaking attacks, thereby enhancing their transferability. The detailed algorithm is summarised in Appendix B.

### v3

@@ Abstract
The integration of new modalities enhances the capabilities of multimodal large language models (MLLMs) but also introduces additional vulnerabilities. In particular, simple visual jailbreaking attacks can manipulate open-source MLLMs more readily than sophisticated textual attacks. However, these underdeveloped attacks exhibit extremely limited cross-model transferability, failing to reliably identify vulnerabilities in closed-source MLLMs. In this work, we analyse the loss landscape of these jailbreaking attacks and find that the generated attacks tend to reside in high-sharpness regions, whose effectiveness is highly sensitive to even minor parameter changes during transfer. To further explain the high-sharpness localisations, we analyse their feature representations in both the intermediate layers and the spectral domain, revealing an improper reliance on narrow layer representations and semantically poor frequency components. Building on this, we propose a Feature Over-Reliance CorrEction (FORCE) method, which guides the attack to explore broader feasible regions across layer features and rescales the influence of frequency features according to their semantic content. By eliminating non-generalizable reliance on both layer and spectral features, our method discovers flattened feasible regions for visual jailbreaking attacks, thereby improving cross-model transferability. Extensive experiments demonstrate that our approach effectively facilitates visual red-teaming evaluations against closed-source MLLMs. Our implementation is released at https://github.com/tmllab/2026_CVPR_FORCE.
@@ 1 Introduction
To meet the growing demand for complex tasks, the capability to process multimodal information has been rapidly integrated into multimodal large language models (MLLMs) [30, 2, 12, 16]. Despite their remarkable performance, the increasing deployment of these models in decision-critical domains has raised societal concerns about their potential risks [31, 11]. Recent red-teaming efforts reveal that, although MLLMs exhibit strong safeguards against textual jailbreaking attacks, they can be easily manipulated through vulnerabilities introduced by newly embedded modalities [32, 6].
Among various attacks, optimisation-based visual jailbreaking attacks are considered one of the most effective for identifying vulnerabilities in MLLMs, as they can reliably bypass the safety guardrails of open-source models with imperceptible perturbations [57, 29, 1]. As illustrated in Figure 1, visual attacks optimised on the source model can effectively exploit its inherent vulnerabilities and elicit harmful responses to malicious instructions, whereas the same requests are refused when paired with a non-adversarial image. Nevertheless, these visual attacks exhibit extremely limited cross-model transferability [35], as the exploited vulnerabilities are specific to the source MLLM and fail to generalise to target MLLMs during transfer. Consequently, such attacks fall short of posing a practical threat to closed-source commercial MLLMs and remain inadequate for real-world red-teaming evaluations.
@@
Motivated by this, we further analyse the feature representations of visual jailbreaking attacks in both the intermediate layers and the spectral domain, uncovering the existence of non-generalizable reliance. Specifically, the feasible regions of visual attacks display distinct characteristics across layers. Closer to the earlier layers, these attacks depend more heavily on model-specific features to mislead MLLMs, resulting in narrower and more fragile feasible regions. Regarding the spectral domain, we observe that as optimisation progresses, high-frequency information exerts increasing influence on attack effectiveness, eventually surpassing low-frequency components that contain richer semantic content. This trend suggests an overemphasis on high-frequency information, making the generated attacks depend on semantically weak features that lack generalisability. Both aspects of improper feature reliance hinder visual jailbreaking attacks from capturing robust representations, which in turn confines them to high-sharpness regions and results in poor transferability.
@@ 2 Related Work
Multimodal Large Language Models. There are two mainstream architectures for integrating new modalities: adapter-based MLLMs [23, 63, 4] and early-fusion MLLMs [62, 48, 39]. Adapter-based MLLMs employ an adapter to project the output of an image encoder, such as CLIP [33], into the embedding space of the large language models (LLMs). On the other hand, early-fusion MLLMs utilise a unified tokeniser to process multimodal information within a shared embedding space. Both designs can leverage the powerful reasoning and understanding capabilities of LLMs to support a wide range of multimodal tasks, with outputs predicted according to the joint conditional distribution of textual and visual information, $p_{\boldsymbol{\theta}}\left(\mathbf{y}\mid\mathbf{x}_{\text{img}},\mathbf{x}_{\text{txt}}\right)$.
@@
Textual Jailbreaking Attack. Jailbreaking attacks arise from the discovery that hand-crafted adversarial prompts can bypass safeguards in LLMs, leading them to answer malicious queries and produce harmful content [38, 25]. To automatically uncover vulnerabilities in LLMs, three types of jailbreaking attack strategies have been rapidly developed. Heuristic-based attacks typically leverage genetic algorithms to modify a prototype corpus until they successfully bypass the safety guardrails [24, 36, 53, 21]. LLM-based attacks utilise the inherent capabilities of LLMs to rewrite malicious queries, obstructing the victim model’s perception [7, 52]. Optimisation-based attacks define an affirmative target output and leverage gradient information to iteratively update the adversarial suffix, ultimately eliciting undesirable responses [64, 50, 20].
@@
Although the aforementioned textual attacks can also manipulate MLLMs, their effectiveness diminishes with the growing strength of textual alignment [41, 5, 34]. In contrast, MLLMs demonstrate relatively weak alignment regarding vulnerabilities associated with new modalities [37, 35], thereby establishing visual jailbreaking attacks as a promising direction for red-teaming evaluations.
@@
Visual Jailbreaking Attack. Visual jailbreaks are typically classified into two categories: generation-based and optimisation-based methods. Generation-based methods either craft image typography to encode malicious textual content [19, 51] or generate harmful images matching the textual semantics [40, 56, 13]. These generated malicious images can mislead MLLMs through the visual modality while simultaneously circumventing textual alignment mechanisms. However, such methods depend on human effort or auxiliary models to produce required visual typography or query–image pairs, making them resource-intensive. More importantly, this type of method lacks the ability to capture the fine-grained vulnerabilities, falling short of reliably manipulating the MLLMs [35].
@@
In contrast, optimisation-based methods, such as the Projected Gradient Descent (PGD) attack [26] and its variants [57, 32, 6, 29], use gradient information to optimise the jailbreaking perturbation $\delta$, thereby reliably exposing model vulnerabilities. In these methods, an affirmative target output of length $S$, such as ‘‘Sure, here is’’, is first defined, and then the loss is calculated as:
@@
Despite achieving near-perfect success in manipulating the source MLLM, optimisation-based methods generate visual attacks with limited transferability to target MLLMs [35]. To thoroughly assess potential risks in closed-source LLMs, this work aims to understand and improve the transferability of optimisation-based visual jailbreaking attacks.
@@ 3.1 Loss Landscape of Visual Jailbreaking Attack
As shown in Figure 1, while optimisation-based visual jailbreaking attacks can easily bypass the safety guardrails of victim MLLMs, their limited transferability to target models constrains their real-world practicality. Inspired by prior research on classification tasks [8, 46], we first investigate the transferability of visual jailbreaking attacks through the geometry of the loss landscape. Throughout this section, we use LLaVA-v1.5-7B [22] as the source MLLM, adopt standard PGD [26] with a step size of $2/255$ and a perturbation budget of $32/255$, and set ‘‘Sure, here is’’ as the optimisation target.
Figure 2: The input (left) and weight (right) loss landscape of the visual jailbreaking attack. The blue and yellow points correspond to successful and failed examples on the source MLLM, respectively.
First, we visualise the input loss landscapes of visual jailbreaking attacks by introducing pixel perturbations in two directions, one aligned with the gradient ascent and the other randomly sampled from a uniform distribution. As observed in Figure 2 (top), the generated visual attacks effectively manipulate the source MLLM to achieve the optimisation objective, as evidenced by the nearly 0 loss at the original point. However, when we inject small pixel perturbations, the loss increases sharply, reflecting that the attack rapidly loses its effectiveness in misleading the model. For instance, even a 0.03 pixel perturbation along the adversarial direction can raise the loss above 0.28, which is sufficient to invalidate the attack. We also introduce weight perturbations to the model parameters to simulate the impact of transfer-induced parameter shifts on attack effectiveness. As depicted in Figure 2 (bottom), we observe that the attack is trapped in a local optimum of the source MLLM, where even a minor weight perturbation of 0.0002 can push it out of the feasible region and render it ineffective. This sharp loss landscape indicates that optimisation-based methods tend to rely on model-specific features, which are sensitive to minor changes and result in unreliable performance when generalised to target models.
@@ 3.2 Features Representations on Different Layers
To disentangle the feature reliance responsible for high-sharpness regions, we conduct a detailed analysis of the intermediate layer representations of generated visual attacks. For a fair comparison, we separately extract each layer’s features from a successful visual jailbreaking attack and a natural image, and then construct interpolated representations using the convex combination $(1-\mu)\cdot f_{\theta}(\text{jail})+\mu\cdot f_{\theta}(\text{nat})$, to exclude inter-layer differences such as parameter norms and activation scales. We also interpolate features between two different visual jailbreaking examples in Appendix A.
@@
However, toward shallower layers, visual jailbreaking attacks exhibit progressively narrower feasible regions in the feature space. As evidenced by Figure 3, in the 11th layer, the attack must retain more than 90% of adversarial features to successfully manipulate the source MLLM, while the introduction of merely 30% of natural features is sufficient to drive the loss sharply beyond 1.2. These observations suggest that in shallower layers, visual jailbreaking attacks exhibit an increasing reliance on model-specific features, manifested as narrower feasible regions. This reliance on non-generalizable early-layer features, in turn, confines the generated attacks to high-sharpness regions of input space, making them unstable when transferred to other models. We also verify the universality of early-layer dependency on different model architectures, as detailed in Appendix B.
@@ 3.3 Influence of Different Frequency Features
In addition to layer-wise features, we also examine the role of spectral information in visual jailbreaking attacks during the optimisation. Specifically, we first apply a Fourier transform to the visual attack and divide the spectrum into ten equal-width frequency bands [18]. Then, we independently mask each frequency band and reconstruct the image via inverse Fourier transform. Finally, we compute the loss of the masked attacks to evaluate their reliance on spectral features.
@@
Nevertheless, as optimisation proceeds, the attack’s effectiveness becomes increasingly dependent on high-frequency components. As shown in Figure 4, at the 350th iteration, the 50–60% and 60–70% spectral features exhibit a more pronounced influence than at the 250th iteration, and removing them causes a greater degradation in attack effectiveness than the lower-frequency 40–50% range. This anomalous trend intensifies with further optimisation. By the 750th iteration, removing the third-highest frequency band alone is sufficient to make the visual jailbreaking attack fail to mislead the source MLLM. This trend indicates that visual jailbreaking attacks tend to increasingly rely on high-frequency features to mislead MLLMs, grounding their success in superficial patterns rather than semantically meaningful content. Such overemphasis on non-generalizable features makes the generated attacks highly model-specific and undermines their transferability across different MLLMs. We also verify the universality of high-frequency dependency on different model architectures, as detailed in Appendix C.
@@
3.4 Feature Over-Reliance CorrEction Method
$d_{l}$ $=\left\|f_{\theta,l}\!\left(\mathbf{x}_{\text{img}}+\delta,\,\mathbf{x}_{\text{txt}}\right)-f_{\theta,l}\!\left(\mathbf{x}_{\text{img}}+\delta+\eta,\,\mathbf{x}_{\text{txt}}\right)\right\|_{2}^{2},$
$\quad\,\,\,l=1,\dots,L.$ (3)
To identify the spectral features with excessive influence, we separately mask $M$ equal-width frequency bands $B_{m}$, and calculate their associated losses $\ell_{m}$, similar to Section 3.3. To restore the natural distribution, where semantic content plays a principal role in model perception, we downscale high-frequency components whenever their influence exceeds the $\beta$-scaled influence of the adjacent low-frequency band:
$w_{m}$ $=\min\!\left(\beta,\tfrac{\ell_{m-1}}{\ell_{m}}\cdot\beta\right),\quad m=1,\ldots,M.$ (7)
$S$ $=\sum\nolimits_{m=1}^{M}(w_{m}\cdot\mathbbm{1}_{B_{m}}).$
Subsequently, we perform an element-wise multiplication of the frequency scaling matrix $S$ with the magnitude spectrum $A$ obtained from the Fourier transform $(A,\Phi)\leftarrow\mathrm{FFT}(\delta)$, and reconstruct the jailbreaking perturbation via the inverse Fourier transform $\delta_{\text{rescaled}}=\mathrm{IFFT}\left((A\odot S)\odot e^{\mathrm{i}\Phi}\right)$. We integrate these two components into a standard PGD algorithm by first rescaling the abnormal frequency bands and then exploring broader layer representations. This design eliminates non-generalizable feature reliance and encourages a flatter loss landscape for the generated visual jailbreaking attacks, thereby enhancing their transferability. The detailed algorithm is summarised in Appendix D.